A sample containing 4.60 g of O2 gas has an initial volume of 19.0 L . What is the final volume, in liters, when each of the following changes occurs in the quantity of the gas at constant pressure and temperature?

A sample of 0.300 mol of O2 is added to the 4.60 g of O2 in the container.

A sample of 1.65 g of O2 is removed.

A sample of 3.20 g of O2 is added to the 4.60 g of O2 gas in the container.
Answer #1

Sol.

As V1 / n1 = V2 / n2  

where

V1 and V2 are initial and final volumes respectively

n1 and n2 are initial and final moles respectively

As V1 = 19.0 L

and n1 = initial Mass of O2 / Molar Mass of O2

= 4.60 / 32 = 0.14375 mol

(a)

A sample of 0.300 mol of O2 is added to 4.60 g of O2 in the container .

So , n2 = 0.14375 + 0.300 = 0.44375 mol

Therefore , V2

= ( V1/n1) × n2

= ( 19.0 / 0.14375 ) × 0.44375

= 58.65 L  

(b)  

A sample of 1.65 g of O2 is removed .

So , n2 = 0.14375 - (1.65/32) = 0.0921875 mol

Therefore ,

V2 = ( V1/n1) × n2

= (19.0/0.14375) × 0.0921875

= 12.18 L

(c)  

A sample of 3.20 g of O2 is added to the 4.60 g of O2 in the container .

So , n2 = 0.14375 + (3.20/32) = 0.24375 mol

Therefore ,

V2 = ( V1/n1) × n2

= ( 19.0 / 0.14375 ) × 0.24375

= 32.22 L